OpenStudy (uscrnamc):

what type of translation takes the graph of (f)x = |x| to the graph g(x) = |4 + x|? A. vertical translation 4 units up B. horizontal translation of 4 unit right c. vertical translation of 4 units down d. horizontal translation of 4 units left

OpenStudy (amorfide):

if I had f(x)=x then f(x+1)=x+1 this is a translation of -1 units parallel to the x axis f(x)+4=x+4 this is a translation of 4 units parallel to the y axis so in your case f(x)=|x| you not have |4+x| so instead of a function of x, we now have 4+x=y so f(y)=|4+x| since y is expressed in terms of x, we can call it a new function g(x) so g(x)=|4+x| so what translation is this?

OpenStudy (amorfide):

a more general term for a translation in x and y directions f(x)=x f(x+a)=x+a, translation -a parallel to x axis f(x)=x f(x)+b=x+b, translation b parallel to y axis